An economy produces three goods: trucks, DVD players, and apples. Quantities and prices per unit for years 2008 and 2009 are shown in the table below 2008 2009 Quantit 13 Price $4,800 $56 $1.00 Quantit Trucks DVD Players Apples Price $7,200 $76 $1.00 1,000 1,000 2008 2009 Nominal GDP $64,184.00 $139,472.00 Average Prices 2008-2009 Trucks DVD Players Apples $6,000.00 S66.00 $1.00 Using the average prices for all three goods, real GDP in 2008 is (Round your response to two decimal places.) Using the average prices for all three goods, real GDP in 2009 is $. (Round your response to two decimal places.) Real GDP from 2008 to 2009 changes by %. (Round your response to two decimal places.) Using the average prices, the GDP deflator for 2008 is (Round your response to two decimal places.) Using the average prices, the GDP deflator for 2009 is. (Round your response to two decimal places.) Using the GDP deflators you calculated above, the rate of inflation for 2008 to 2009 is | %. (Round your response to two decimal places.) Enter your answer in each of the answer boxes
Answer #1

Real GDP in 2008 = 13*6000 + 14*66 + 1000*1 = $79924

Real GDP in 2009 = 19*6000 + 22*66 + 1000*1 = $116452

Real GDP changes from 2008 to 2009 = (116452 - 79924)/79924

Real GDP changes from 2008 to 2009 = 45.7%

GDP deflator for 2008 = ((13*4800 + 14*56 + 1000*1)/79924)*100

GDP deflator for 2008 = 80.31

GDP deflator for 2009 = ((19*7200 + 22*76 + 1000*1)/116452)*100

GDP deflator for 2009 = 119.77

Rate of inflation from 2008 to 2009 = (119.77 - 80.31)/80.31

Rate of inflation from 2008 to 2009 = 49.13%

So, the answer is as follows:

Using the average prices for all three goods, real GDP in 2008 is $ . (Round your response to two decimal places.) 79924 Usin